On average, 1 in 75000 adults has a certain genetic disorder.
- (a)Use a suitable approximating distribution to find the probability that, in a random sample of 10000 people, at least 1 has the genetic disorder.[3]
- (b)In a random sample of n people, where n is large, the probability that no-one has the genetic disorder is more than 0.9. Find the largest possible value of n.[4]
Mark scheme
Marking points from the official mark scheme (May/June 2021). The mark or mark type is shown in brackets.
- [M1] For Poisson approximation parameter `λ = (1/75000) * 10000 = 10000/75000 = 10/75 = 2/15`. SOI. Allow `Po(0.133)`.
- [M1] For calculating `P(X >= 1) = 1 - P(X=0) = 1 - e^(-λ)`. Allow incorrect λ, allow one end error.
- [A1] Final probability `= 0.125 (3 sf)`. SC Partially unsupported final answer: `Po(2/15)` stated B1 then unsupported `0.125` B1. SC Use of Binomial (`0.1248`) B1 only. Use of Normal scores M0.
- [B1] For Poisson parameter `λ = n/75000`.
- [M1] For inequality `e^(-n/75000) > 0.9`. Allow '='.
- [M1] For taking natural logarithm on both sides: `-n/75000 > ln 0.9` [leading to `n < 7902.04`].
- [A1] Largest possible value of n is `7902`. CWO. Must be an integer.
Write your answer and get it marked
On Oracle Prep you can write or photograph your own answer and have it marked by AI against this mark scheme, with feedback on what you missed.
Sign up free →